OpenStudy (anonymous):

x-1/x^2-1 - x/x-3 + x^2/ x^2-2x-3 it would help ALOT!

jhonyy9 (jhonyy9):

x2 -1 = (x-1)(x+1) x2 -2x -3 = (x+1)(x-3)

jhonyy9 (jhonyy9):

after this check the denominator common

OpenStudy (anonymous):

\[\frac{1}{x+1} -\frac{x}{x-3} + \frac{x^2}{(x-3)(x+1)}\]